OpenStudy (anonymous):

There are two right circular cylinders that have the same height, 45 centimeters. The radius of the smaller cylinder is 22 centimeters and the radius of the larger cylinder is 6 times greater than that of the smaller cylinder. What is the ratio of the volume of the larger cylinder to the volume of the smaller cylinder? a. 6:1 b. 12:1 c. 36:1 d. 150:1 Will someone please explain to me how to do this? Thanks in advanced!!! :)

OpenStudy (zepp):

Alright, what's the radius of the larger cylinder?

OpenStudy (zepp):

Radius of the large is 6 times greater than the smaller one, therefore radius is \(22*6=132\)

OpenStudy (anonymous):

I got that. But i seem to be missing something.

OpenStudy (zepp):

Now find the volume of the small cylinder; Height = 45cm Radius = 22cm \(\large V_{Cyl}=A_{Base}*\text{Height}\) \(\large V=\pi r^2*h\) \(\large V=\pi*22^2*45\) \(\large V=21780\pi\) Is that what you got?

OpenStudy (zepp):

Following the same formula, find the volume of the big cylinder Height = 45cm Radius = 132cm \(\large V=\pi r^2*h\\V=\pi*132^2*45\\V=784080\pi\)

OpenStudy (anonymous):

not at all

OpenStudy (zepp):

"What is the ratio of the volume of the larger cylinder to the volume of the smaller cylinder?" \(\large V_{Large}=784080\pi\) \(\large V_{Small}=21780\pi\) \[\large \text{Ratio}=\frac{784080\pi}{21780\pi}=\frac{36}{1}\]

OpenStudy (zepp):

Can you show me your steps? So I can point out where are your mistakes and what you should remember the next time you'll encounter this kind of problem :)

OpenStudy (anonymous):

My steps endedwith 22*6. i was lost. but i am reworking now to try to figure it out. thanks.

OpenStudy (zepp):

You are welcome, if you have any problem, just message me like you did earlier :)

OpenStudy (anonymous):

Will do! Thanks again!!!